OpenStudy (anonymous):

can anyone help me with precal? Find tan θ if sec θ = square root of ten divided by three and sin θ < 0. please help?

OpenStudy (adajiamcneal):

tan(theta)=+- sqrt((sec(theta))^2-1 )= +- sqrt(10/9 -1)=+-1/3 hence tan(theta)=-1/3 as sin θ < 0

OpenStudy (anonymous):

what?